OpenStudy (anonymous):

Solve for x. -x + 4 = 32 x = -36 x = -28 x = 28 x = 36

OpenStudy (anonymous):

-x + 4 = 32 - 4 = -4 Subtract 4 from each side. -x = 28 ----------- -x = 28 -x = -x Divide -x from both sides ( In other words divide -1 from both sides). x =-28 Answer